Two identical capacitors are connected in parallel to an ac generator that has a frequency of 523 Hz and produces a voltage of 24 V. The current in the circuit is 0.14 A. What is the capacitance of each capacitor?
Answer #1

We know that

Z = V/i

Z = 24/0.14 = 171.43 ohm

Now

Z = Xc = 1/w*Ceq

Ceq = 1/(w*Xc)

Ceq = 1/(2*pi*523*171.43)

Ceq = 1.775*10^-6 F

for two identical capacitors connected in series,

Ceq = C1 + C2 = 2*C

C = capacitance of each capacitor

C = Ceq/2

C = 0.887*10^-6 C

C = 0.89 uF